OpenStudy (anonymous):

8x + 7y = 39 4x – 14y = –68

OpenStudy (anonymous):

Use elimination. Multiply the bottom by -2 would be easiest. 8x +7y =39 -8x+28y=136 then add them. 0x=35y=175 175 divided by 35 is 5, so y=5 then plug 5 in for y anywhere 8x + 7(5) = 39 8x +35 = 39 8x= 4 x=1/2 , y=5